OpenStudy (anonymous):

HELPP ME PLEASE !Choose the correct description of the graph of the compound inequality 2x + 5 > 9 and 5x less than or greater to 30 A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right. A number line with an open circle on 2, a closed circle on 6, and shading in between. A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right. A number line with a closed circle on 2, an open circle on 6, and shading in between.h

OpenStudy (anonymous):

can you solve the inequalities ?

OpenStudy (anonymous):

No im really bad at it

OpenStudy (anonymous):

they are solved just like equations except for the fact, that when dividing/multiplying by a negative number, the inequality sign changes

OpenStudy (anonymous):

can you solve equations ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

2x + 5 > 9.......pretend this is an equation....can you solve for x ?

OpenStudy (anonymous):

its x>2

OpenStudy (anonymous):

now do the other inequality the same way

OpenStudy (anonymous):

I believe the other inequality is supposed to be less then or equal

OpenStudy (anonymous):

x<=6

OpenStudy (anonymous):

so we have 2x greater then 2 and x less then or equal to 6. Quick lesson... If it is greater then, shading goes to the right, if it is less then, shading goes to the left. Only use a closed circle if there is an equal sign, if there is not an equal sign involved, use an open circle.

OpenStudy (anonymous):

2x+5>9 2x>9-5 2x>4 x>2 ...(1) 5x <>30 x< .6 ,x not equal to 6 ...(2) from (1) & (2) \[x \in \left( 2,6 \right)\cup \left( 6,\infty \right)\]

OpenStudy (anonymous):

oops...x greater then 2

OpenStudy (anonymous):

so is the answer a

OpenStudy (anonymous):

no......try again

OpenStudy (anonymous):

kelliegirl33?

OpenStudy (anonymous):

C ?

OpenStudy (anonymous):

x is greater then 2, and less then or equal to 6

OpenStudy (anonymous):